OpenStudy (anonymous):

You drive in a straight line at 24.0 m/s for 8.0 miles, then at 30.0 m/s for another 8.0 miles. What is the average velocity?

OpenStudy (lgbasallote):

average velocity = total distance/ total time first time = distance/speed = 8/24 = 0.33 s second time = distance/speed = 8/30 = 0.27 s therefore \[V_{ave} = \frac{8 + 8}{0.33 + 0.27}\] got it?

OpenStudy (anonymous):

we first gotta convert 8 miles in metres... then we can proceed in the same manner as igbasallote did !!